OpenStudy (anonymous):

One option in a roulette game is to bet $8 on red. There are 18 red compartments, 18 black compartments, and two compartments that are neither red not black. If the ball lands on red, you get to keep the $8 you paid to play the game and you are awarded $8. If the ball lands elsewhere, you are awarded nothing and the $8 bet is collected. What is the expected value for playing roulette if you bet $8 on red? Round your answer to the nearest cent.

OpenStudy (anonymous):

expected value: the amount you win times the probability you win it which of course includes the possibility of losing. you bet $8 on red and you win with probability \[\frac{18}{38}\] and lose $8 with probability \(\frac{20}{38}\) (that is how they get you!)

OpenStudy (anonymous):

so compute \[8\times \frac{18}{38}-8\times \frac{20}{38}\]

OpenStudy (anonymous):

the answer should be in fraction ?

OpenStudy (anonymous):

\[\frac{8\times 18-8\times 20}{38}=\frac{-16}{38}=-\frac{8}{19}=-.421\] rounded

OpenStudy (anonymous):

I got -8/19

OpenStudy (anonymous):

oh it says "round your answer to the nearest cent" so i guess you need \(-.42\)

OpenStudy (anonymous):

it turns out that no matter what you bet on roulette, blocks, even, odd, red, black, whatever the expected value is always the same, \(-\frac{1}{18}\) for every dollar bet

OpenStudy (anonymous):

Oh I get it thank you so much you're a life saver

OpenStudy (anonymous):

typo there it is \(-\frac{1}{19}\) for every dollar

OpenStudy (anonymous):

yw my pleasure

OpenStudy (anonymous):

on line class?

OpenStudy (anonymous):

Yeah :( That's the hard part about it

OpenStudy (anonymous):

Hate Online Classes

OpenStudy (anonymous):

yeah a drag, except that you can always find other resources on line to help

OpenStudy (anonymous):

Oh well I guess I just have to keep up with it until i graduate Thank's for the heads up

OpenStudy (anonymous):

yw